# Decomposing Global Feature Effects Based on Feature Interactions

### v2

@@ D.2 Proof of Theorem 8
$\mathds{P}\biggl{(}\mathcal{R}(\widehat{f}_{\mathcal{D}_{n}^{\Pi}},\tilde{\mathbf{x}}_{j},\mathcal{X})>\epsilon\biggr{)}\to 0$
@@
$\quad\mathds{P}\biggl{(}\mathcal{R}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})\leq z_{1-\alpha}\biggr{)}$ (41)
$\leq\mathds{P}\biggl{(}\mathcal{R}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})\leq z_{1-\alpha},z_{1-\alpha}\geq c\biggr{)}+\mathds{P}\biggl{(}\mathcal{R}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})\leq z_{1-\alpha},z_{1-\alpha}<c\biggr{)}$
$\leq\mathds{P}\biggl{(}z_{1-\alpha}\geq c\biggr{)}+\mathds{P}\biggl{(}\mathcal{R}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})<c\biggr{)}$ (42)
@@
Appendix E Further Empirical Validation of PINT
E.1 Example for Failure of PINT
@@
E.2 Further Empirical Validation of PINT
@@
Appendix F GADGET-SD Results for COMPAS
@@
The effect plots for the four resulting regions are shown in Figures 20 and 21. GADGET based on SD finds the same first split as for PD. The second split is also executed according to the number of prior crimes, but the split value is lower than for PD (at $2.5$ instead of $4.5$). The total interaction-related heterogeneity reduction ($R^{2}_{Tot}=0.87$) are also similar. Note that Shapley values explain the difference between the actual and average prediction. Hence, SD plots are centered, while Figure 11 shows the uncentered regional PD plots.
@@
Appendix G Further Details on Higher Dimensional Settings
@@
Figure 22: Regional PD plots for remove, credit and edu of the spam data set after applying GADGET when only interacting features are selected for $S$ and $Z$ according to the pre-filtering steps. The grey areas indicate the $95\%$ interaction-related heterogeneity interval as defined in Appendix B.4.
Appendix H Rashomon Effect
@@
Appendix I Marginal- and Conditional-based Feature Effect Methods
@@
While marginal-based approaches are usually simple to define and efficiently computable, they are known to extrapolate into unseen regions of the feature space when features are highly correlated, which might lead to misleading interpretations as illustrated in Section 4.6 for PD plots (Hooker, 2007; Molnar et al., 2022). Conditional-based approaches solve the extrapolation problem by considering the true underlying data distribution (Aas et al., 2021). However, they require estimating the conditional data distribution, which is challenging, particularly in high-dimensional settings. The estimation influences the quality of the derived interpretation (Aas et al., 2021; Sundararajan and Najmi, 2020; Janzing et al., 2020). While both approaches come with different merits and weaknesses, the derived interpretation often differs as well. For example, PD plots visualize the influence of a feature of interest on the prediction function irrespective of the underlying feature correlations. In contrast, the M plot accounts for the correlation structure and thus also takes into account the influence of the features correlated with the feature of interest. Therefore, the two approaches answer different questions: Marginal-based approaches primarily offer insights into the inner workings of the ML model without explicitly considering the underlying data correlations, making them particularly useful for model audit and debugging. Conditional-based approaches are used to derive interpretations that reflect the given data distribution and thus are preferable for questions related to model inference (Chen et al., 2020; Freiesleben et al., 2022; Watson, 2022).
To conclude, Chen et al. (2020) argue that the marginal-based approach is not generally wrong and can be useful if we want to derive explanations that are true to the model, while the conditional-based approach should be used when we want to extract interpretations that are true to the data. The framework that we introduce in this paper works with both, marginal-based (e.g., PD plots as illustrated in Section 4.3) as well as conditional-based (e.g., ALE as shown in Section 4.4) approaches. Hence, the user needs to decide on a suitable feature effect method depending on their desired interpretation.

### v3

@@ E.2 Proof of Theorem 8
$\mathds{P}\biggl{(}\mathcal{R}_{j}(\widehat{f}_{\mathcal{D}_{n}^{\Pi}},\tilde{\mathbf{x}}_{j},\mathcal{X})>\epsilon\biggr{)}\to 0$
@@
$\quad\mathds{P}\biggl{(}\mathcal{R}_{j}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})\leq z_{1-\alpha}\biggr{)}$ (43)
$\leq\mathds{P}\biggl{(}\mathcal{R}_{j}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})\leq z_{1-\alpha},z_{1-\alpha}\geq c\biggr{)}+\mathds{P}\biggl{(}\mathcal{R}_{j}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})\leq z_{1-\alpha},z_{1-\alpha}<c\biggr{)}$
$\leq\mathds{P}\biggl{(}z_{1-\alpha}\geq c\biggr{)}+\mathds{P}\biggl{(}\mathcal{R}_{j}(\widehat{f}_{\mathcal{D}_{n}},\tilde{\mathbf{x}}_{j},\mathcal{X})<c\biggr{)}$ (44)
@@
Appendix F Further Empirical Validation of PINT
F.1 Example for Failure of PINT
@@
F.2 Further Empirical Validation of PINT
@@
Appendix G GADGET-SD Results for COMPAS
@@
The effect plots for the four resulting regions are shown in Figures 21 and 22. GADGET based on SD finds the same first split as for PD. The second split is also executed according to the number of prior crimes, but the split value is lower than for PD (at $2.5$ instead of $4.5$). The total interaction-related heterogeneity reduction ($R^{2}_{Tot}=0.87$) is also similar. Note that Shapley values explain the difference between the actual and average prediction. Hence, SD plots are centered, while Figure 12 shows the uncentered regional PD plots.
@@
Appendix H Further Details on Higher Dimensional Settings
@@
Figure 23: Regional PD plots for remove, credit and edu of the spam data set after applying GADGET when only interacting features are selected for $S$ and $Z$ according to the pre-filtering steps. The grey areas indicate the $95\%$ interaction-related heterogeneity interval as defined in Appendix C.4.
Appendix I Rashomon Effect
@@
Appendix J Marginal- and Conditional-based Feature Effect Methods
@@
While marginal-based approaches are usually simple to define and efficiently computable, they are known to extrapolate into unseen regions of the feature space when features are highly correlated, which might lead to misleading interpretations as illustrated in Section 4.6 for PD plots (Hooker, 2007; Molnar et al., 2022). Conditional-based approaches solve the extrapolation problem by considering the true underlying data distribution (Aas et al., 2021). However, they require estimating the conditional data distribution, which is challenging, particularly in high-dimensional settings. The estimation influences the quality of the derived interpretation (Aas et al., 2021; Sundararajan and Najmi, 2020; Janzing et al., 2020). While both approaches come with different merits and weaknesses, the derived interpretation often differs as well. For example, PD plots visualize the influence of a feature of interest on the prediction function irrespective of the underlying feature correlations. In contrast, the M plot accounts for the correlation structure and thus also takes into account the influence of the features correlated with the feature of interest. Therefore, the two approaches answer different questions: Marginal-based approaches primarily offer insights into the inner workings of the ML model without explicitly considering the underlying data correlations, making them particularly useful for model audit and debugging. Conditional-based approaches are used to derive interpretations that reflect the given data distribution and thus are preferable for questions related to model inference (Chen et al., 2020; Freiesleben et al., 2024; Watson, 2022).
To conclude, Chen et al. (2020) argue that the marginal-based approach is not generally wrong and can be useful if we want to derive explanations that are true to the model, while the conditional-based approach should be used when we want to extract interpretations that are true to the data. The framework that we introduce in this paper works with both marginal-based (e.g., PD plots as illustrated in Section 4.3) as well as conditional-based (e.g., ALE as shown in Section 4.4) approaches. Hence, the user needs to decide on a suitable feature effect method depending on their desired interpretation.